# Exploring Robustness of LLMs to Sociodemographically-Conditioned Paraphrasing

### v1

@@ 4.1 Performance on Paraphrased Sets
Table 3: Performance comparison of paraphrases generated with Semantic Information Augmented re-Generation using embedding similarity of original-paraphrases sentences as the augmented information in the prompt (0-shot QA evaluation)
@@ 4.2 Semantically Augmented Re-Generation
When we examine the semantic closeness of the paraphrased sets to the original contexts (Figure 5), we observe that the Ambiguous gender category and the Young age group are the lower-performing sets in terms of both prediction accuracy and average semantic similarity scores when compared to their counterparts and the original sets. This highlights a correlation between the quality of the generated paraphrases for these sets and the performance on the QA task. Table 3 presents a detailed performance comparison of paraphrases generated using semantic information augmented re-generation, which incorporates embedding similarity of paraphrased sentences to the original context as the augmented information in the prompt. We perform this assessment on zero-shot QA evaluation. The embeddings are derived using the Sentence Transformers architecture with the pre-trained weights of "paraphrase-distilroberta-base-v2". The model is specifically trained on a large corpus of paraphrase data, making it highly effective for tasks involving paraphrase identification.
@@ 5.1 Human evaluation of paraphrases
We conducted a detailed analysis of paraphrase predictions to assess the impact of stylistic and semantic changes on their validity across male and female sets. The misclassifications were categorized into two primary types: incorrect paraphrases and hallucinations. Incorrect paraphrases were those that did not preserve the original meaning, while hallucinations involved the addition of information not present in the source text. We analyzed the misclassification that are common (shared between male and female sets), male-only, and female-only (Table 4). The common set, which encompasses errors found in both male and female samples, shows an invalid paraphrasing rate of approximately 25%, including both incorrect renditions and hallucinations. This rate is higher than that observed in the male-only set but lower than in the female-only set. Such a pattern suggests a slight correlation between the overall misclassification rates for the male and female sets, possibly hinting that the model adds more artifacts to the paraphrases for the female context. Generally, these artifacts are related to verbs of niceness/politeness which alter the meaning of the original sentence.
@@
5.2 Correlation with Perplexity
Perplexity is a measure of a model’s uncertainty in predicting the next word in a sequence. In the context of linguistic variation, given that the content of the sentence is the same, perplexity can also be interpreted as how comfortable is the model with predicting that style. The perplexity scores reported in Table 1 for LLAMA2 and GPT-2 across different demographic categories reveal significant insights into model behavior.
Notably, we see an inverse relationship between perplexity and model performance; as perplexity increases, we observe a corresponding decrease in model accuracy. Figure 6 shows this correlation. We consider GPT-2 as an unbiased ’third’ model to calculate the set perplexity because the language is generated by the LLM itself. For GPT-2, results with the 0-shot setup demonstrate an inverse correlation between perplexity and model accuracy and higher perplexity scores directly translate to lower accuracy. This trend is also evident, albeit not so strongly, in the 2-shot setting. This possibly indicates that when the model is allowed to learn the language nuances in a low data setup, the role of language complexity (perplexity) plays a smaller part.
@@ 5.3 Effects of Language Diversity
The exploration of language diversity’s impact on QA performance uncovers notable trends in model robustness across various demographic sets. Specifically, our analysis indicates that the Young and Gender Ambiguous demographic sets exhibit the least robustness in QA tasks when compared to the rest of the categories, such as Male, Female, Middle Age, and Old. This pattern of performance sheds light on potential biases and gaps within the training datasets utilized for developing the LLMs, suggesting that the Young and Gender Ambiguous sets might be underrepresented.
Furthermore, the reduction in performance seems to stem from the nature of the language employed in the paraphrasing for these demographics, which tends to be more expressive, with a higher perplexity, and less formal compared to other groups. It suggests that the model’s ability to understand and process information may be significantly influenced by the formality of the language. This implies that the commonsense reasoning capacity of a model could vary with the formal structure of the language, with less formally written language presenting a greater challenge for accurate comprehension of the social or commonsense context. Such findings highlight a crucial consideration for the development of more robust QA models. We need to account for and adapt to the linguistic diversity inherent in human language, particularly when it concerns informal or expressive language styles.
@@ 5.4 Semantic Differences in Paraphrasing
Misclassifications. We explore the role of semantic differences in paraphrases and their impact on misclassifications in QA tasks, Figure 7. We measure the cosine similarity between the original and paraphrase embeddings obtained through the ’paraphrase-distilroberta’ model. Upon manual verification, it was observed that out of the common misclassified samples in the Male and Female set, only about 20% could be accounted for by incorrect/hallucinated paraphrasing. The rest of the misclassifications were likely due to linguistic changes.
@@
Explainability of paraphrases with XSBERT. We examine the performance of LLMs on paraphrased sets in comparison to the semantic attributions between original and paraphrased sentences. We rely on explainability methods that enhance our understanding of model decisions, which are obtained with the integration of XSBERT Moeller et al. (2023) and DistilRoBERTa Sanh et al. (2019). XSBERT is an adaptation of the Siamese network architecture applied to sentence transformers. It simultaneously processes paired inputs and maps them onto a scalar output, which is represented as model attribution scores and attribution errors for an input sentence pair. Traditional methods like cosine similarity are effective for determining the likeness of the sentences overall but do not account for token-wise attributions. In contrast, XSBERT utilizes integrated Jacobians which provide a richer and matrix-based attribution that measures the strength of relationships and the specific features contributing to model decisions. The trends seen in Figure 8 suggest that the explainability of paraphrasing plays a crucial role in the model’s assessment of the context.
@@ 6 Conclusion
The integration of demographic-specific paraphrasing significantly impacts the performance of QA models as evidenced by our experiments with the LLaMA2 and MISTRAL models. Our analysis revealed that demographic nuances alter the model’s ability to accurately interpret and answer questions. For instance, paraphrases tailored to younger demographics, which often incorporate slang and informal expressions, tend to decrease model accuracy. This suggests a potential gap in the model’s training data or its ability to handle less formal language, which is commonly used by younger individuals. On the other hand, paraphrases associated with Old, Male, and Middle Age demographics, which typically feature more formal language, showed better compatibility with the model’s predictive capabilities, indicating that LLMs are more familiar with less expressive language and a more formal tone. This differential impact underscores the need for models that are robust across varied linguistic styles and demographic contexts. Through deeper analysis of model reliability, we explored the quantifiable aspects of linguistic variations that affect the reasoning capabilities. We find that informal and expressive language associated with Young and Ambiguous sets is higher in complexity and harder to reason with. The disparity is also visible in the type of question being asked as evidenced by the ATOMIC analysis (Appendix F). This highlights a potential scope of improvement for the LLMs to be able to perform tasks equally well in different language expressions.
@@ 8 Acknowledgements
Appendix A Optimal LLaMA Hyperparameters
@@
Prompt style LexHub alignment %($M<F$) Cumulative scores
Controlled 72% Ambiguous: 1.33
Female: 235.90
Male: -248.36
Style guided 64% Ambiguous: 12.49
Female: 112.67
Male: -86.97
Table 8: Paraphrase alignment with LexHub for controlled vs. style-guided paraphrasing with LLAMA2. The cumulative scores show the raw aggregated scores for the paraphrased samples.
Appendix B Evaluation Formula
@@
Appendix C Demographic Artefacts in Paraphrases
@@
Appendix D Semantic Clusters for Paraphrases
@@
Appendix E Sentiment Alignment with LexHub
@@
Appendix F Performance on ATOMIC questions
ATOMIC knowledge is categorized into social interactions, physical events, and mental states, with dimensions such as causes (xIntent, xNeed), effects (xEffect, oEffect), and attributes (xAttr, oReact). To label the SIQA development set with ATOMIC labels, we manually map the social situations and behaviors described in SIQA to the corresponding dimensions and categories found in the ATOMIC dataset. The analysis of ATOMIC label interpretation across demographics within the SIQA dataset reveals distinct patterns in how LLM predictions are affected when different groups process social scenarios, shown in Figure 11. For example, younger individuals often struggle with xNeed, which is reflective of their ability to describe their emotional states. Middle-aged paraphrasing improves the interpretation of social cues in xIntent and xReact, and the understanding of motivations and emotional responses is improved. The impact of paraphrasing is not uniform across all ATOMIC categories. For instance, xNeed (needs before an action) and oWant (desires post-event) often see greater fluctuations in accuracy. This could be because the subtleties of needs and desires are sensitive to linguistic differences. The paraphrased datasets generally show close accuracy in recognizing xAttr (attributes of individuals) and xReact (reactions to social situations), implying that describing attributes and predicting reactions is not extremely dependent on language style.

### v2

@@ 4.2 Encoder Models
Table 4: Performance comparison of paraphrases generated with Semantic Information Augmented Re-generation using embedding similarity of original paraphrases as the augmented information in the prompt (0-shot QA evaluation)
Figure 6: Performance with paraphrased sets contrasted with the semantic similarity of the paraphrased set with the SIQA validation set. The colored bars represent the demographic set accuracy while the gray bars show their respective cosine similarity score.
@@ 4.3 Semantically Augmented Re-generation
When we examine the semantic closeness of the paraphrased sets to the original contexts (Figure 6), we observe that the Ambiguous gender category and the Young age group are the lower-performing sets in terms of both prediction accuracy and average semantic similarity scores when compared to their counterparts and the original sets. This highlights a correlation between the quality of the generated paraphrases for these sets and the performance on the QA task. Table 4 presents a detailed performance comparison of paraphrases generated using semantic information augmented re-generation, which incorporates embedding similarity of paraphrased sentences to the original context as the augmented information in the prompt. We perform this assessment on zero-shot QA evaluation. The embeddings are derived using the Sentence Transformers architecture with the pre-trained weights of paraphrase-distilroberta-base-v2 Reimers and Gurevych (2019). The model has been trained on a large corpus of paraphrase data, making it suitable for tasks involving paraphrase identification.
For the LLAMA2 model, results indicated nuanced differences across demographic groups when comparing the effects of one regeneration versus three. For instance, the Male group saw an improvement increase from a negligible decrease after one regeneration to a 0.96% improvement after three regenerations. The Female group demonstrated a similar progressive improvement, from 1.18% after one regeneration to 1.69% after three, showcasing a clear benefit of multiple regeneration cycles in refining paraphrases. In contrast, the Young group’s performance continued to decline, with -0.62% after one regeneration and -0.71% after three, indicating that semantic reinforcement may not be enough to enhance understanding of the Young style.
The MISTRAL model displayed somewhat different dynamics. The Male group initially experienced a decrease of 0.23% after one regeneration but improved by 0.25% after three regenerations. The Female group also showed significant improvement over time, with a slight initial drop turning into a 0.88% gain after three regenerations. This pattern was also visible albeit less pronounced in the Old group which already shows a high 0-shot accuracy. We see slight but steady improvements over the initial performance, highlighting a slow but beneficial adaptation to regeneration.
When comparing the two models, LLAMA2 tends to show more significant performance fluctuations across multiple regenerations, indicating a potential for higher gains but also highlighting its sensitivity to the regeneration process. Meanwhile, MISTRAL tends to show a steadier progression or recovery in performance, especially notable in groups that initially experience a drop. Overall, the analysis shows that generally, multiple regenerations can enhance paraphrase quality.
@@ 5.1 Human Evaluation of Paraphrases
We conducted a detailed analysis of paraphrase predictions to assess the impact of stylistic and semantic changes on their validity across male and female sets. The misclassifications were categorized into two primary types: incorrect paraphrases and hallucinations. Incorrect paraphrases were those that did not preserve the original meaning, while hallucinations involved the addition of information not present in the source text.
We analyzed the misclassification that are common (shared between male and female sets), male-only, and female-only (Table 5). The common set, which encompasses errors found in both male and female samples, shows an invalid paraphrasing rate of approximately 25%, including both incorrect renditions and hallucinations. This rate is higher than that observed in the male-only set but lower than in the female-only set. Such a pattern suggests a slight correlation between the overall misclassification rates for the male and female sets, possibly hinting that the model adds more artifacts to the paraphrases for the female context. Generally, these artifacts are related to verbs of niceness/politeness which alter the meaning of the original sentence.
@@ 5.2 Effects of Language Diversity
The exploration of language diversity’s impact on QA performance uncovers notable trends in model robustness across various demographic sets. Specifically, our analysis indicates that the Young and Gender Ambiguous demographic sets exhibit the least robustness in QA tasks when compared to the rest of the categories, such as Male, Female, Middle Age, and Old (Figure 2). Furthermore, the reduction in performance seems to stem from the nature of the language employed in the paraphrasing for these demographics, which tends to be more expressive, with a higher perplexity, and less formal compared to other groups (Table 9 in Appendix). It suggests that the formality of the language may significantly influence the model’s ability to process information. This implies that the commonsense reasoning capacity of a model could vary with the formal structure of the language, with less formally written language presenting a greater challenge for accurate comprehension of the social or commonsense context. Such findings highlight a crucial consideration for developing more robust QA models, taking into account the linguistic diversity inherent in human language such as informality or expressiveness.
@@ 5.3 Semantic Differences in Paraphrasing
We investigate the role of semantic differences in paraphrases and their impact on misclassifications in QA tasks (Figure 7). We measure the cosine similarity between the original and paraphrase embeddings obtained from the paraphrase-distilroberta model Reimers and Gurevych (2019). Upon manual verification, it was observed that out of the common misclassified samples in the Male and Female set, only about 20% could be accounted for by incorrect/hallucinated paraphrasing. The rest of the misclassifications were likely due to linguistic changes.
@@ 6 Conclusion
Our analysis revealed that demographic nuances alter the model’s ability to accurately interpret and answer questions. Paraphrases tailored to younger demographics, which often use informal expressions, tend to decrease model accuracy. This suggests a potential gap in the model’s training data or its ability to handle informal language. On the other hand, paraphrases associated with Old, Male, and Middle Age demographics, which typically feature more formal language, showed better compatibility with the model’s predictive capabilities, indicating that LLMs are more familiar with less expressive language and a more formal tone. Our findings underscore the need for more robust models across the linguistic styles of different demographics.
@@ 8 Acknowledgements
Appendix A Explainability of Paraphrases with XSBERT.
We examine the performance of LLMs on paraphrased sets in comparison to the semantic attributions between original and paraphrased sentences. We rely on explainability methods that enhance our understanding of model decisions, which are obtained with the integration of XSBERT Moeller et al. (2023) and DistilRoBERTa Sanh et al. (2019). XSBERT is an adaptation of the Siamese network architecture applied to sentence transformers. It simultaneously processes paired inputs and maps them onto a scalar output, which is represented as model attribution scores and attribution errors for an input sentence pair. Traditional methods like cosine similarity are effective for determining the likeness of the sentences overall but do not account for token-wise attributions. In contrast, XSBERT utilizes integrated Jacobians which provide a richer and matrix-based attribution that measures the strength of relationships and the specific features contributing to model decisions. The trends seen in Figure 8 suggest that the explainability of paraphrasing plays a crucial role in the model’s assessment of the context.
Appendix B Two-shot Baselines
In Table 6, we observe some of the existing 2-shot baselines taken from BIG-Bench with popular LLMs. The reported values are an average of three runs of the SIQA test set and paraphrased sets. Compared to the existing baselines, LLAMA2 performs significantly better. It is also worth noting that with our proposed answer selection approach using log-likelihoods, we see an improvement over the officially reported baseline scores with LLAMA2.
Model Accuracy Config
Random 33.3 -
GPT-3 Small 33.7 2 shot
GPT-3 XL 35.5 2 shot
GPT-3 200B 46.8 2 shot
PaLM 64B 58.0 2 shot
LLAMA2 (original) 50.4 0 shot
LLAMA2 (ours) 57.8 0 shot
MISTRAL 66.3 0 shot
Table 6: SIQA 2-shot Baselines from BIG-Bench.
Appendix C Optimal LLaMA Hyperparameters
@@
Appendix D Evaluation Formula
@@
Appendix E Demographic Artefacts in Paraphrases
@@
Appendix F Semantic Clusters for Paraphrases
@@
Appendix G Sentiment Alignment with LexHub
@@
Appendix H Performance on ATOMIC questions
ATOMIC knowledge is categorized into social interactions, physical events, and mental states, with dimensions such as causes (xIntent, xNeed), effects (xEffect, oEffect), and attributes (xAttr, oReact). To label the SIQA development set with ATOMIC labels, we manually map the social situations and behaviors described in SIQA to the corresponding dimensions and categories found in the ATOMIC dataset. The analysis of ATOMIC label interpretation across demographics within the SIQA dataset reveals distinct patterns in how LLM predictions are affected when different groups process social scenarios, shown in Figure 11. For example, younger individuals often struggle with xNeed, which is reflective of their ability to describe their emotional states. Middle-aged paraphrasing improves the interpretation of social cues in xIntent and xReact. The impact of paraphrasing is not uniform across all ATOMIC categories. For instance, xNeed (needs before an action) and oWant (desires post-event) often see greater fluctuations in accuracy. This could be because the subtleties of needs and desires are sensitive to linguistic differences. The paraphrased datasets generally show close accuracy in recognizing xAttr (attributes of individuals) and xReact (reactions to social situations), implying that describing attributes and predicting reactions is not extremely dependent on language style.